A square loop of wire 10.0 cm on a side is in a 1.25 T magnetic field B. What are the maximum and minimum values of flux that can pass through the loop?

1
Expert's answer
2022-04-27T13:34:09-0400
"\Phi_{min}=0"

"\Phi_{max}=BA=Ba^2\\
=\rm 1.25\: T\times (0.1\: m\times 0.1\; m)=0.0125\: Wb"